find the slope and y-intercept of the line 7x+4y=24

Found 2 solutions by friesr, Alan3354:
Answer by friesr(113)

The form is y=mx+b
m is the slope
7x+4y = 24
4y = -7x + 24
y = -7/4x + 6
Slope is -7/4
set Y = to 0
0 = -7/4x + 6
-6 = -7/4x
x = -6 * -4/7
x = 24/7
(24/7, 0)

Answer by Alan3354(69443)

find the slope and y-intercept of the line 7x+4y=24
-------
Solve for y
y = (-7/4)x + 6
slope = -7/4, the coefficient of x
y-int = 6, the other term